40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 29 18:10:13 2009, Yoichi, Update, PSL, ISS Bad OPLTF1.png
    Reply  Fri Jan 30 19:38:57 2009, Kakeru, Update, PSL, ISS Bad 
Message ID: 1260     Entry time: Thu Jan 29 18:10:13 2009     Reply to this: 1262
Author: Yoichi 
Type: Update 
Category: PSL 
Subject: ISS Bad 
Kakeru, Yoichi

As we noted before, the ISS is unstable. You can see the laser power oscillation around 3Hz.
We took the open-loop transfer function of the ISS around the lower UGF.
The phase margin is almost non-existent.
It was measured with the ISS gain slider at 2dB (usually it was set to 7dB).
So if we increase it by 3dB, it is guaranteed to be unstable.

The higher UGF has also a small phase margin (about 12deg.).
With the ISS gain slider at 2dB, the upper UGF is too low, i.e. the UGF is located at the beginning of the 1/f region.
So we if we make the lower UGF stable by lowering the gain, the upper UGF becomes unstable.

We took out the ISS box from the PSL table.
Kakeru and Peter are now trying to modify the filter circuit to give more phase margin at the lower UGF.
Attachment 1: OPLTF1.png  147 kB  | Hide | Hide all
OPLTF1.png
ELOG V3.1.3-